OpenStudy (anonymous):

brason and lily are working together on solving math problems. for every one problem that brason completes, lily completes 7. if the number of problems solved by lily is fifty more than two times the number of problems solved by brason,how many problems has lily solved?

OpenStudy (anonymous):

Wow. Why would Lily bother working with Brason, I wonder? :) In any case, I am supposed to be working on the number of problems Lily solved, not the group dynamics. :) It seems your problem can be solved by writing two linear equations and solving it. Since Lily solves 7 problems for every one Brason solves, the number of problems Lily solved (x) is 7 times the number of problems Brason solved (y), or x=7y Since Lily solves 50 more than twice the problems Brason solves, we can represent it with 2y+50=x. Therefore, by the Transitive property, 2y+50=7y, 50=5y and y=10 Therefore Lily solved 70 problems. :)

OpenStudy (anonymous):

L = 7B L = 50 + 2B -> 7B = 50 + 2B => 5B = 50 => B = 10 ( Brandon solves 10 problems) Thus Lily solves 7 * 10 = 70 problems